### v3

@@ B. Different Groups of States Join the Union: War Participants
For religion, we use the number of (Southern) Baptist churches per 1,000 capita as a proxy for a more pro-slavery religious environment, and the number of Friends (Quaker) churches per 1,000 capita as a proxy for an antislavery religious environment. Prior researchers have noted that among the Protestant denominations, (Southern) Baptists were generally more supportive of slave ownership. This followed the split of the Baptist Church in 1845, which produced the Southern Baptist Convention after national mission boards refused to appoint slaveholders as missionaries. This has been interpreted by historians as an indication of (Southern) Baptist support for the legitimacy of slave holding prior to the Civil War (Mathews, 1977; Noll, 2006). Although the split was less clear in the border states—where Northern and Southern affiliations coexisted—Baptists in these regions still tended to oppose abolitionism (Harlow, 2014). Other Protestant churches (e.g., Methodists and Presbyterians) had more diverse views regarding slavery (Wigger, 2001; Snay, 1993).
@@
As shown in Appendix Table S9, including these variables and their interactions with peer exposure does not alter our main findings concerning peer influence among cadets from Slave States nor the interaction of the peer effect with slave population share. Among the newly added controls, the antislavery vote share (Lincoln) is positively associated with the probability of joining the Union, consistent with expectations.
These results also suggest that Slave Population Share and Cadet Slave Ownership are more powerful proxies for the most relevant economic, social, and political considerations than these other proxies.
@@ 3.4 Shared Experience and Peer Influence
A significant share—26.1%—of cadets in our dataset participated in the Mexican-American War (1846–1848), offering a valuable context to examine how shared experiences shape peer influence. For each cadet, we compute the proportion of peers who also fought in the war, distinguishing between those from slave and Free States.
@@
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Share Free-State peers join Mexican-American War: The proportion of Free-State peers participating in Mexican-American war with themselves. If the cadet did not participate in the Mexican-American War, the proportion is 0. Share Slave-State peers joining the Mexican-American war: The proportion of Slave-State peers participating in Mexican-American War with themselves. If the cadet did not participate in the Mexican-American War, the proportion is 0. The control variables consist of Age in 1860, Class Rank, state level Slave Population Share (sd) and Cohort. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
@@ 3.5 Additional Analyses
First, we divide the sample into periods before and after 1850. Although the issue of slavery was highly contentious throughout the period we study, it was particularly intense in the 1850s and nearly led to the collapse of the United States. In an effort to avert civil war, Congress passed the Compromise of 1850, which included the Fugitive Slave Law that obligated law enforcement nationwide to assist in capturing alleged runaway slaves, and many in the north resisted its enforcement (McPherson, 2003). Against this backdrop, we use 1850 as a dividing point to assess whether peer influence experienced notable changes before and after that year. As illustrated in Supplemental Appendix Table S11, peer influence persisted over time, with similar effects on either side of this divide.
@@
We know the post-war outcomes of many who joined an army, allowing us to investigate how their decisions during this critical historical juncture influenced their ex post life results, such as military rank and survival probability. We do not presume that this had any influence on their allegiances, but does trace some consequences of their decisions.
@@ 4 Discussion
Our study highlights the significant role of peer exposure in shaping West Point cadets’ allegiance choices during the American Civil War, especially for those from Slave States who faced a tension between nationalism and sectionalism among other tensions. Importantly, we find that the strength of peer influence was modulated by the economic interests of cadets’ home states: in regions where the slave economy was overwhelmingly dominant, the influence of peers was less pronounced. This suggests that when economic stakes were particularly high and slavery more prevalent, cadets were less susceptible to peer-driven shifts in loyalty. This finding is also consistent with the fact that states with heavier slave shares were the first to secede and the level of debate in states was at least partly reflective of slave shares (Crofts, 2014). Thus, the places in which cadets could make choices without seeming disloyal were states with relatively lower slave population shares.
Although our analysis is a case study, it is a useful proof of concept. The decisions that West Point cadets made during the U.S. Civil War parallel those that arise in many divided societies, where individuals must choose between between highly-polarized opposing sides during critical moments. We have shown that in such a context, peer interactions can still play a decisive and significant role.
Our study also provides a possibility for identifying such peer effects in other high-stakes and high-polarization settings. This sort of peer influence is challenging to quantify and identify causally through anecdotal evidence alone. It is also not enough to identify correlation in people’s behaviors, given that friendships are endogenous and people are subject to common (potentially unobserved) factors (Angrist, 2014). West Point’s environment—characterized by structured, close-knit interactions among cadets from diverse geographic backgrounds—provided us an opportunity to analyze the influence of quasi-random peer exposure, and its interplay with background allegiances. By using variation in a cadet’s peers’ state-of-origin for identification, we have something that is exogenous to other key things influencing a given cadet’s decision. Similar techniques can be used in other high-stakes contexts.
That said, our ability to definitively trace the mechanisms behind peer influence is limited by the nature of our historical data. Our specification that examines cadets’ choices based on the composition of their cohort by the states of origin of their peers enables us to causally identify peer influence. Nonetheless, it does not provide insight into why peers influence cadets’ decisions.
@@
Theoretically, however, we are still left with the fact that peer effects could have operated in at least three ways: (1) by fostering communication and persuasion, whereby cadets from Slave States may have been convinced to support the Union cause through exposure to pro-Union peers, (2) by building friendship bonds that influenced cadets’ decisions to fight alongside their peers, or (3) by shifting perceptions of the likely outcome of the war, where cadets from Free States could have led their peers to believe that the Union was more likely to prevail. Our data, however, do not allow us to disentangle these, or other mechanisms, and all could have operated together (which would be consistent with some evidence from personal letters and diaries (McPherson, 1997)). Further investigation of peer influence on allegiances in modern contexts, where additional data are more readily available, could offer additional insights into how such mechanisms operate.
@@ A.3 Summary Statistics
Slave Pop. Share (County-level) 24.269 18.178 0.446 71.269 244
@@
Slave Pop. Share (County-level) 0.000 0.000 0 0.011 433
@@ A. Cohort level bootstrap
Share Free-State peers (sd) 0.080∗∗∗ 0.058∗∗∗ 0.054∗∗∗ -0.003 -0.005 -0.006
(0.020) (0.014) (0.014) (0.015) (0.015) (0.015)
Controls N Y Y N Y Y
State FEs N N Y N N Y
Dependent var. mean 0.361 0.361 0.361 0.919 0.919 0.919
Observations 388 388 388 540 540 540
R-squared 0.028 0.234 0.287 0.000 0.017 0.030
@@ B. Cohort level cluster
Share Free-State peers (sd) 0.080∗∗∗ 0.058∗∗∗ 0.054∗∗∗ -0.003 -0.005 -0.006
(0.020) (0.013) (0.019) (0.012) (0.012) (0.013)
Controls N Y Y N Y Y
State FEs N N Y N N Y
Dependent var. mean 0.361 0.361 0.361 0.919 0.919 0.919
Observations 388 388 388 540 540 540
R-squared 0.028 0.234 0.287 0.000 0.017 0.030
Note. Panel A reports the re-estimation of the baseline results using robust standard errors clustered at the cohort level. Panel B reports the corresponding re-estimates using bootstrapped standard errors clustered at the cohort level with 400 resampling iterations. (This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. The control variables consist of Age in 1860, Class Rank, state level Slave Population Share (sd), and Cohort. Columns (1)-(3) focus on cadets from Slave States and Columns (4)-(6) on those from Free States. ***$p<0.01$, **$p<0.05$, *$p<0.1$.)
@@ B.3 Home State Economic Proxies and War Allegiances
Share Free-State Peers (sd) 0.058∗∗∗ 0.058∗∗∗ 0.054∗∗ 0.053∗∗ -0.005 -0.006 -0.005 -0.005
@@ B.4 Religion and Voting
(1) (2) (3) (4) (5)
Share Free-State Peers (sd) 0.053∗∗ 0.054∗∗ 0.055∗∗∗ 0.056∗∗ 0.054∗∗
(0.021) (0.021) (0.021) (0.022) (0.022)
Slave Share (County) $\times$ Share Free-State Peers (sd) -0.036 -0.034 -0.036 -0.047∗∗ -0.041∗∗
(0.018) (0.018) (0.020) (0.021) (0.020)
Slave Share (County) -0.166∗∗∗ -0.164∗∗∗ -0.159∗∗∗ -0.142∗∗∗ -0.161∗∗∗
(0.033) (0.033) (0.034) (0.034) (0.034)
Baptists $\times$ Share Free-State Peers (sd) 0.005
Baptists -0.040
FriendQuaker $\times$ Share Free-State Peers (sd) 0.004
FriendQuaker 0.022
Modern Republican $\times$ Share Free-State Peers (sd) -0.013
Modern Republican 0.091∗∗
Southern Democrat $\times$ Share Free-State Peers (sd) -0.003
Southern Democrat -0.021
Controls Y Y Y Y Y Y
State FEs Y Y Y Y Y Y
Dependent var. mean 0.358 0.358 0.358 0.330 0.330
Observations 352 352 352 327 327
R-squared 0.341 0.347 0.342 0.349 0.337
Note. Our sample for this table is restricted to cadets from Slave States. All regressions include standardized versions of Slave share (county), Baptists, FriendsQuaker, Modern Republican, and Southern Democrat. We exclude 36 observations with missing county-level information. The control variables consist of Age in 1860, Class Rank, and Cohort. Columns (4)–(5) have smaller sample sizes because Washington, DC did not have complete voting records prior to 1861. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.

### v4

@@ B. Different Groups of States Join the Union: War Participants
As a proxy for underlying religious attitudes, we follow Fogel (1994) by classifying denominations into two broad groups: evangelical churches (Methodists, Presbyterians, and Baptists) and liturgical churches (Catholics, Lutherans, and Episcopalians), with the former tending to be associated with pro-slavery constituencies, and the latter tending to be associated with anti-slavery constituencies. We use the number of churches of the given group per 1000 population as the measure.
@@
As shown in Appendix Table S9, among the additional variables, evangelical churches per capita is negatively associated with the probability of joining the Union, while the antislavery vote share (Lincoln) is positively associated with it. We also see that including these variables and their interactions with peer exposure does not alter our main findings concerning peer influence among cadets from Slave States nor the interaction of the peer effect with slave population share. We note that the absence of a significant interaction between peer influence and religion may reflect the imprecision of our proxy for underlying religious attitudes, limited statistical power, or the fact that several denominations were internally divided, making it difficult to fully capture the role of religion in allegiance decisions (Wigger, 2001; Snay, 1993).
@@ 3.4 Shared Experience and Peer Influence
A significant share—30.5%—of cadets in our dataset participated in the Mexican-American War (1846–1848), offering a valuable context to examine how shared experiences shape peer influence. For each cadet, we compute the proportion of peers who also fought in the war, distinguishing between those from slave and Free States.
@@
Note. Our sample for this table is restricted to cadets from Slave States whose cohorts fall between 1820 and 1845. It reports the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union army. Share Free-State peers join Mexican-American War: The proportion of Free-State peers participating in Mexican-American war with a cadets. If the cadet did not participate in the Mexican-American War, the proportion is 0. Share Slave-State peers joining the Mexican-American war: The proportion of Slave-State peers participating in Mexican-American War with themselves. If the cadet did not participate in the Mexican-American War, the proportion is 0. The control variables consist of Age in 1860, Class Rank, state level Slave Population Share (sd) and Cohort. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
@@ 3.5 Additional Analyses
First, we divide the sample into periods before and after 1850. Although the issue of slavery was highly contentious throughout the period we study, it was particularly intense in the 1850s. In an effort to avert civil war, Congress passed the Compromise of 1850, which included the Fugitive Slave Law that obligated law enforcement nationwide to assist in capturing alleged runaway slaves, and many in the north resisted its enforcement (McPherson, 2003). Against this backdrop, we use 1850 as a dividing point to assess whether peer influence experienced notable changes before and after that year. As illustrated in Supplemental Appendix Table S11, peer influence persisted over time, with similar effects on either side of this divide.
@@
We observe the post-war outcomes of many who joined an army, allowing us to investigate how their decisions during this critical historical juncture influenced their ex post life results, such as military rank and survival probability. We do not presume that this had any influence on their allegiances, but it does trace some consequences of their decisions.
@@ 4 Discussion
Our study highlights the significant role of peer exposure in shaping West Point cadets’ allegiance choices during the American Civil War, especially for those from Slave States who faced a tension between nationalism and sectionalism among other tensions. We find that the strength of peer influence was modulated by the economic interests of cadets’ home states: in regions where the slave economy was large, the influence of peers was less pronounced. This suggests that when economic stakes were particularly high and slavery more prevalent, cadets were less susceptible to peer-driven decisions. This finding is also consistent with the fact that states with larger slave shares were the first to secede and the level of debate in states was at least partly reflective of slave shares (Crofts, 2014).
Although our analysis is a case study, it is a useful proof of concept. The decisions that West Point cadets made during the U.S. Civil War parallel those that arise in many divided societies, where individuals must choose between between polarized factions during critical moments. We have shown that in such a context, peer interactions can still play a decisive and significant role.
Our study also provides a possibility for identifying such peer effects in other high-stakes and polarized settings. This sort of peer influence is challenging to quantify and identify causally through anecdotal evidence alone. It is also not enough to identify correlation in people’s behaviors, given that friendships are endogenous and people are subject to common (potentially unobserved) factors (Angrist, 2014). West Point’s environment—characterized by structured, close-knit interactions among cadets from diverse geographic backgrounds—provided us an opportunity to analyze the influence of quasi-random peer exposure, and its interplay with background allegiances. By using variation in a cadet’s peers’ state-of-origin for identification, we have something that is exogenous to other key things that influenced a given cadet’s decision. Similar techniques can be used in other high-stakes contexts.
@@
We are still left with the fact that peer effects could have operated in (at least) three ways: (1) by fostering communication and persuasion, whereby cadets from Slave States may have been convinced to support the Union cause through exposure to pro-Union peers, (2) by building friendship bonds that influenced cadets’ decisions to fight alongside their peers, or (3) by shifting perceptions of the likely outcome of the war, where cadets from Free States could have led their peers to believe that the Union was more likely to prevail. Our data do not allow us to disentangle these, or other mechanisms, and all could have operated together (which would be consistent with some evidence from personal letters and diaries (McPherson, 1997)). Further investigation of peer influence on allegiances in modern contexts, where additional data are more readily available, could offer additional insights into how such mechanisms operate.
@@ A.3 Summary Statistics
Slave Pop. Share (County-level) 24.269 18.718 0.446 71.269 244
@@
Slave Pop. Share (County-level) 0.000 0.000 0 0.012 433
@@ A. Cohort level bootstrap
Share Free-State Peers (sd) 0.080∗∗∗ 0.058∗∗∗ 0.054∗∗∗ 0.056∗∗∗ -0.003 -0.005 -0.006 (0.015)
(0.020) (0.014) (0.014) (0.017) (0.015) (0.015) (0.015) (0.015)
Controls N Y Y Y N Y Y Y
State FEs N N Y Y N N Y Y
Dependent var. mean 0.361 0.361 0.361 0.355 0.919 0.919 0.919 0.930
ICC 0.000 0.000 0.000 0.000 0.028 0.016 0.015 0.025
Observations 388 388 388 324 540 540 540 456
R-squared 0.028 0.234 0.287 0.363 0.000 0.017 0.030 0.067
@@ B. Cohort level cluster
Share Free-State Peers (sd) 0.080∗∗∗ 0.058∗∗∗ 0.054∗∗∗ 0.056∗∗∗ -0.003 -0.005 -0.006 -0.005
(0.020) (0.013) (0.019) (0.020) (0.012) (0.012) (0.013) (0.012)
Controls N Y Y Y N Y Y Y
State FEs N N Y Y N N Y Y
Dependent var. mean 0.361 0.361 0.361 0.355 0.919 0.919 0.919 0.930
ICC 0.000 0.000 0.000 0.000 0.028 0.016 0.015 0.025
Observations 388 388 388 324 540 540 540 456
R-squared 0.028 0.234 0.287 0.363 0.000 0.017 0.030 0.067
Note. Panel A reports the re-estimation of the baseline results using robust standard errors clustered at the cohort level. Panel B reports the corresponding re-estimates using bootstrapped standard errors clustered at the cohort level with 400 resampling iterations. (This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. The control variables consist of Age in 1860, Class Rank, state level Slave Population Share (sd), and Cohort. In addition, columns (4) and (8) further control for cadet slave ownership as well as family background characteristics, including whether the cadet’s father, mother, or wife was from a free state. Columns (1)-(4) focus on cadets from Slave States and Columns (5)-(8) on those from Free States. ***$p<0.01$, **$p<0.05$, *$p<0.1$.)
@@ B.3 Home State Economic Proxies and War Allegiances
Share Free-State Peers (sd) 0.058∗∗∗ 0.058∗∗∗ 0.054∗∗∗ 0.053∗∗∗ -0.005 -0.006 -0.005 -0.005
@@ B.4 Considering Religion and Voting
Slave States Border States
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
Share Free-State Peers (sd) 0.053∗∗ 0.053∗∗ 0.054∗∗ 0.056∗∗ 0.054∗∗ 0.068∗∗∗ 0.066∗∗∗ 0.068∗∗∗ 0.073∗∗ 0.074∗∗
(0.021) (0.021) (0.021) (0.022) (0.022) (0.025) (0.025) (0.025) (0.029) (0.029)
Slave Share (County) $\times$ Share Free-State Peers (sd) -0.036 -0.035 -0.043∗∗ -0.047∗∗ -0.041∗∗ -0.027 -0.026 -0.028 -0.046 -0.037
(0.018) (0.019) (0.020) (0.021) (0.020) (0.023) (0.023) (0.025) (0.026) (0.023)
Slave Share (County) -0.166∗∗∗ -0.170∗∗∗ -0.166∗∗∗ -0.142∗∗∗ -0.161∗∗∗ -0.187∗∗∗ -0.190∗∗∗ -0.185∗∗∗ -0.163∗∗∗ -0.183∗∗∗
(0.033) (0.033) (0.034) (0.034) (0.034) (0.036) (0.036) (0.036) (0.036) (0.034)
Evangelical $\times$ Share Free-State Peers (sd) -0.007 0.013
(0.018) (0.024)
Evangelical -0.039 -0.029
(0.024) (0.033)
Liturgical $\times$ Share Free-State Peers (sd) -0.022 -0.003
(0.024) (0.026)
Liturgical 0.011 0.010
(0.036) (0.045)
Republican $\times$ Share Free-State Peers (sd) -0.013 -0.017
(0.018) (0.022)
Republican 0.091∗∗ 0.090
(0.043) (0.053)
Southern Democrat $\times$ Share Free-State Peers (sd) -0.003 0.003
(0.023) (0.029)
Southern Democrat -0.021 -0.020
(0.045) (0.046)
Controls Y Y Y Y Y Y Y Y Y Y
State FEs Y Y Y Y Y Y Y Y Y Y
Dependent var. mean 0.358 0.358 0.358 0.330 0.330 0.471 0.471 0.471 0.443 0.443
Observations 352 352 352 327 327 244 244 244 219 219
R-squared 0.341 0.346 0.343 0.349 0.337 0.295 0.298 0.295 0.318 0.308
Note. Our sample for this table is restricted to cadets from Slave States. For religion, we use the number of churches per 1,000 population as a proxy for a religious environment. The control variables consist of Age in 1860, Class Rank, and Cohort. Columns (4)–(5) and (9)–(10) have smaller sample sizes because Washington, DC did not have complete voting records prior to 1861. Columns (1)-(5) focus on cadets from Slave States and Columns (6)-(10) on those from Border States. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.